OpenStudy (anonymous):

10% off 400 $ would be 360$ ?

OpenStudy (amistre64):

1% is a division by 100 1% = 400/100 10% is just 10 of those 1%s

OpenStudy (amistre64):

400 - 10% is?

OpenStudy (anonymous):

360.

jigglypuff314 (jigglypuff314):

you are correct :)

OpenStudy (amistre64):

or we could just calculate it out right:\[\frac{360-400}{400}=?~-0.10\]

OpenStudy (anonymous):

Thank you for your time.

OpenStudy (amistre64):

youre welcome